OpenStudy (mcoban):

A laser produces 20.0 mW of red light. In 1 hour, the laser emits 2.29 X 10^20 photons. What is the wavelength of the laser?

OpenStudy (anonymous):

Wavelengh is \[6.31\times 10^{-7}\]

OpenStudy (anonymous):

Do you need my solution?

OpenStudy (anonymous):

yes

OpenStudy (mcoban):

of course I do, please

OpenStudy (anonymous):

The power of laser is \[2\times 10^{-3} Joule/seconds\] so in 1 hour or 3600 seconds we get 72 joules. \[E=nhc/\] we know h, c and n and E, so we get wavelengh.

OpenStudy (mcoban):

sorry but what is n ?

OpenStudy (anonymous):

you are welcomed. n is the number of photons and it has said that it is \[2.29\times 10^{20}\]

OpenStudy (mcoban):

Thank you so much

OpenStudy (anonymous):

It is an honor to help. Do you study molecular spectroscopy?

OpenStudy (mcoban):

I study Quantum-Mechanical Model of Atom